29.2%(w/w) HCL stock solution has a density of 1.25 g/ml.What is the volume of stock solution in ml. reqd. to prepare a 200 ml. solution of 0.4M HCL?

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
29.2 wt/wt solution means 29.2g HCl in 100gm solution 
that means 29.2g HCl in 100/1.25 =80ml. That means 29.2/35.5 =.8 mole in 80 ml so now the solution is 10M 
we need to make .4M solution and total of 200ml. 
so that means a total of .4*.2 =.08 mol of HCl must be there
to bring that lets sat x litter of the previous solution is taken then x*(10)= 0.08 mol so x=.008  l or  ml and rest 192 ml water or other thig